OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 4 and 9. What is the length of the altitude? (4 points) 4 6 9 13

OpenStudy (raden):

x^2 = 4 * 9 solve for x

OpenStudy (anonymous):

what does * stand for?

OpenStudy (raden):

multiply

OpenStudy (anonymous):

Is it x=18

OpenStudy (raden):

x^2 = 4 * 9 x^2 = 36 x = sqrt(36) = ....

OpenStudy (anonymous):

Ohhhh i divided insted of sqrt

OpenStudy (raden):

lol... try again

OpenStudy (anonymous):

I got x=6

OpenStudy (raden):

yep... correct

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (raden):

welcome :)